# Exact Expressive Power of Transformers with Padding

### v1

@@ Appendix B $\mathsf{TC}^{d}$ Circuit Evaluation Problem
We refer to the special case where $\mathsf{C}=\mathsf{TC}^{d}$ as the $\mathsf{TC}^{d}$ circuit evaluation problem. This corresponds to threshold circuit evaluation where the depth of the circuit is constrained to be $O(\log^{d}\lvert x\rvert)$.
@@
We further define wide-$\mathsf{TC}^{d}\subseteq\mathsf{TC}^{d}$ as the class of circuit families $\{C_{n}\}_{n=0}^{\infty}$ such that there exists some $c$ such that, for large $n$, the depth of $C_{n}$ is at most $c\log^{d}n$ and, crucially, the size is at least $n^{c}$. That is, the size (and hence the width) of the circuit is large relative to its depth.
We define the corresponding wide-$\mathsf{TC}^{d}$ circuit evaluation problem as a variant of circuit evaluation problem with $\mathsf{C}=\text{wide-}\mathsf{TC}^{d}$.
B.1 Hardness of $\mathsf{TC}^{d}$ Circuit Evaluation
@@
For $d\geq 1$, $\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions.
@@
Given any $L\in\mathsf{L}\text{-uniform}\ \mathsf{TC}^{d}$, there exists a log-space Turing machine $T_{L}$ that constructs a circuit family $\{C_{n}\}_{n=0}^{\infty}$ that recognizes $L$. We can construct an $\mathsf{L}$-reduction from $L$ to the $\mathsf{TC}^{d}$ circuit evaluation problem as follows. Given an input $x$ whose membership in $L$ we would like to check, the reduction first copies $x$ to the output. It then uses the log-space Turing machine $T_{L}$ to build the circuit $C_{\lvert x\rvert}$ and output it in the above serialized format. We thus have a log-space reduction from $x$ to $(x,\langle C_{\lvert x\rvert}\rangle)$. We conclude that $\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions. ∎
@@
For $d\geq 1$, wide-$\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions.
@@
As in the proof of Lemma 5, we are given $L$ such that there exists a log-space Turing machine $T_{L}$ that constructs $\{C_{n}\}_{n=0}^{\infty}$ recognizing $L$, where the depth of each $C_{n}$ is at most $c\log^{d}n$ for some $c$ and large enough $n$. We can create a modified log-space Turing machine $T^{\prime}_{L}$ that builds $\{C^{\prime}_{n}\}_{n=0}^{\infty}$ that still recognizes $L$ in depth $c\log^{d}n$ but is padded with dummy gates so that it has size at least $n^{c}$: we do this by keeping a counter for the number of gates and wires produced and outputting dummy ones until $n^{c}$ is exceeded. We then follow the rest of the proof of Lemma 5 with $T^{\prime}_{L}$ instead of $T_{L}$. ∎
B.2 Solving $\mathsf{TC}^{d}$ Circuit Evaluation with Transformers
@@
We adapt the proof of Merrill and Sabharwal [2023a, Theorem 3], which sketches how log-depth (unlooped) transformers can implement the $\mathsf{TC}^{0}$ circuit evaluation problem. We construct a looped transformer that will “attempt” to evaluate every gate: if its arguments have already been computed, the gate will return $\{0,1\}$, and, if not, it will return undefined ($\bot$).
[…]
@@
For any $d\geq 0$, the wide-$\mathsf{TC}^{d}$ circuit evaluation problem is in $\mathsf{mAHAT}^{d}_{0}$, and hence in $\mathsf{AHAT}^{d}_{1}$.
@@
We are given input $(x,\langle C_{n}\rangle)$, where $C_{n}$ comes from some $\mathsf{TC}^{d}$ circuit family $\{C_{n}\}_{n=0}^{\infty}$ with depth at most $c\log^{d}n$ and size at least $n^{c}$, where $c$ is a constant specific to the circuit family. If we unroll the transformer from Lemma 7 to depth $c\log^{d}n$, we can solve wide-$\mathsf{TC}^{d}$ circuit evaluation problem for large enough $n$ (w.l.o.g. we can solve small-$n$ examples via table lookup).
@@
Finally, we conclude via Proposition 1 that this mixed-masked transformer can be converted to a corresponding causally masked padded transformer, placing the problem also in $\mathsf{AHAT}^{d}_{1}$. ∎
Appendix C Uniformity Ceiling: Sanity Checking Our Results
We recount some simple but counterintuitive facts obtained while verifying the consistency of our results. In general, for classes $\mathsf{A},\mathsf{B}$ such that $\mathsf{A}\subseteq\mathsf{B}$, $\mathsf{B}$-uniformity often leads to larger classes of languages than $\mathsf{A}$-uniformity, as we have more resources to construct a circuit. However, as the following proposition shows, this is not always the case.
For the purposes of this result, we assume that $\mathsf{A}$-uniform $\mathsf{XC}$ means there exists $f\in\mathsf{A}$ such that $f$ maps $1^{n}$ to an $\mathsf{XC}$ circuit that operates over inputs of size $n$. This is not satisfied by some stronger notions of uniformity like $\mathsf{DLOGTIME}$ uniformity, but it is satisfied for $\mathsf{L}$ and $\mathsf{NL}$ uniformity.
Proposition 2 (Uniformity ceiling).
Consider classes $\mathsf{A},\mathsf{B}$ and a circuit class $\mathsf{XC}$ such that:
@@ Appendix C Uniformity Ceiling: Sanity Checking Our Results
    $\mathsf{XC}$ circuit evaluation is in $\mathsf{A}$-uniform $\mathsf{XC}$.
Then $\mathsf{B}$-uniformity does not strengthen $\mathsf{A}$-uniformity, i.e., $\mathsf{A}\textrm{-uniform}\;\mathsf{XC}=\mathsf{B}\textrm{-uniform}\;\mathsf{XC}$.
@@
We will show that any language $L\in\mathsf{B}\textrm{-uniform}\;\mathsf{XC}$ is also in $\mathsf{A}\textrm{-uniform}\;\mathsf{XC}$. By definition of $\mathsf{B}\textrm{-uniform}\;\mathsf{XC}$, there exists a procedure $f\in\mathsf{B}$ that maps any input $\texttt{1}^{n}$ to an $\mathsf{XC}$ circuit $C^{f}_{n}$ that checks membership in $L$ over all strings $w$ of size $n$, i.e., $f(\texttt{1}^{n})=C^{f}_{n}$. By the first precondition of the proposition, $f$ is also in $\mathsf{A}\textrm{-uniform}\;\mathsf{XC}$. This means there exists a procedure $g\in\mathsf{A}$ that maps any input $\texttt{1}^{n}$ to an $\mathsf{XC}$ circuit $C^{g}_{n}$ that does what $f$ does, i.e., on input $\texttt{1}^{n}$ constructs the $\mathsf{XC}$ circuit $C^{f}_{n}$. Thus, we have $g(\texttt{1}^{n})=C^{g}_{n}$ and $C^{g}_{n}(\texttt{1}^{n})=C^{f}_{n}$. Furthermore, by the second precondition of the proposition, there exists a procedure $h\in\mathsf{A}$ that maps any input $(\texttt{1}^{n},\texttt{1}^{m})$ to an $\mathsf{XC}$ circuit $C^{h}_{n,m}$ that evaluates any $\mathsf{XC}$ circuit $C_{n}$ of size $m$ over any input $w$ of size $n$, i.e., $h(\texttt{1}^{n},\texttt{1}^{m})=C^{h}_{n,m}$ and $C^{h}_{n,m}(C_{n},w)=C_{n}(w)$.
Now we are ready to construct an $\mathsf{A}$-uniform $\mathsf{XC}$ circuit for recognizing $L$. Using currying notation from functional programming, [^5] Currying is the technique of translating a function that takes multiple arguments into a sequence of families of functions, each taking a single argument, cf. https://en.wikipedia.org/wiki/Currying. As an example, currying turns a two-argument function $f:(x,y)\mapsto z$ into a family of single-argument functions $g:x\mapsto(y\mapsto z)$, i.e., $g(x)$ is a single-argument function that maps $y$ to $z$ (in a manner that depends on $x$). this circuit is $C_{L}=h(\texttt{1}^{n},\texttt{1}^{m})(g(\texttt{1}^{n})(\texttt{1}^{n}))$ where $n=\lvert w\rvert$ and $m=\lvert g(\texttt{1}^{n})\rvert$. Here we use currying notation to specify that the first input of the two-input circuit $C^{h}_{n,m}$ generated by $h$ is fixed as $g(\texttt{1}^{n})(\texttt{1}^{n})$, turning it into a single-input function. Similarly, $g(\texttt{1}^{n})(\texttt{1}^{n})$ denotes first applying $g$ to $\texttt{1}^{n}$ in order to construct a circuit $g(\texttt{1}^{n})=C^{g}_{n}$, and then passing $\texttt{1}^{n}$ as input to this constructed circuit $C^{g}_{n}$.
By design, $C_{L}$ takes as input a string $w$ and computes $C_{L}(w)=h(\texttt{1}^{n},\texttt{1}^{m})(g(\texttt{1}^{n})(\texttt{1}^{n}),w)$.
It is easy to see that $C_{L}$ can be constructed by a procedure in $\mathsf{A}$—after all, $g$ and $h$ are in $\mathsf{A}$, $\texttt{1}^{n}$ can be easily computed from $w$ by converting all 0s to 1s, and similarly $\texttt{1}^{m}$ can be computed from the output $g(\texttt{1}^{n})$ of applying procedure $g$ (which is in $\mathsf{A}$) on $\texttt{1}^{n}$. Further, $C_{L}$ is a circuit in $\mathsf{XC}$ as it is a fixed composition of the $\mathsf{XC}$ circuits that functions $h$ and $g$ output. All that remains to be shown is that $C_{L}$ recognizes $L$. To this end, we note:
$C_{L}(w)$ $=h(\texttt{1}^{n},\texttt{1}^{m})(g(\texttt{1}^{n})(\texttt{1}^{n}),w)$
$=C^{h}_{n,m}(g(\texttt{1}^{n})(\texttt{1}^{n}),w)$
$=C^{h}_{n,m}(C^{g}_{n}(\texttt{1}^{n}),w)$
@@
$=C^{f}_{n}(w).$
By design, $C_{f}(w)=1$ if and only if $w\in L$. Thus, $C_{L}$ recognizes $L$. ∎
Applying this result in the case of $\mathsf{L}$ and $\mathsf{NL}$ as classes $\mathsf{A}$ and $\mathsf{B}$, respectively, we obtain:
Corollary 3.1.
For any circuit class $\mathsf{XC}$ containing $\mathsf{NL}$, $\mathsf{L}\textrm{-uniform}\;\mathsf{XC}=\mathsf{NL}\textrm{-uniform}\;\mathsf{XC}$.

### v2

@@ Appendix D Wide-$\mathsf{TC}^{d}$ Circuit Evaluation Problem
For example, the case where $\mathsf{C}=\mathsf{P/poly}$ yields the generic circuit value problem, which is known to be $\mathsf{P}$-complete. We focus here to the case of $\mathsf{C}=\mathsf{TC}^{d}$, i.e., the $\mathsf{TC}^{d}$ circuit evaluation problem. It is somewhat intuitive that this problem is hard for the class $\mathsf{A}\textrm{-uniform}\ \mathsf{TC}^{d}$ as long as $\mathsf{A}$ is strong enough to build circuits; we formalize this later in Lemma˜11. However, the $\mathsf{TC}^{d}$ circuit evaluation problem is not necessarily in the class $\mathsf{TC}^{d}$. To see this, suppose to the contrary that there exists an $\mathsf{A}\textrm{-uniform}\ \mathsf{TC}^{d}$ circuit family $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ that solves the $\mathsf{TC}^{d}$ circuit evaluation problem. Then, for large enough $n$, each circuit $C_{n}\in\mathcal{C}$ has depth upper bounded by $c\log^{d}n$ for a fixed $c>0$. This $c$ being fixed is problematic—if one were to try to evaluate on string $x$ a circuit $C^{\prime}_{n}$ from a $\mathsf{TC}^{d}$ circuit family that has depth $c^{\prime}\log^{d}n$ where $c^{\prime}>c$ (that is, invoke circuit evaluation for input $(x,\langle C^{\prime}_{n}\rangle)$), the intuitive approach would require a circuit whose depth is larger than that of $C_{n}$.
To address this, we now formalize a constrained version of the $\mathsf{TC}^{d}$ circuit evaluation problem that is, in fact, within the class $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$. We achieve this via the wide-$\mathsf{TC}^{d}$ circuit evaluation problem, where the class wide-$\mathsf{TC}^{d}$ is defined as follows:
Definition 16 (Wide Circuits).
@@
Let wide-$\mathsf{TC}^{d}\subseteq\mathsf{TC}^{d}$ be the class of circuit families $\{C_{n}\}_{n=0}^{\infty}$ such that there exists some $c$ such that, for large $n$, the depth of $C_{n}$ is at most $c\log^{d}n$ and, crucially, the size is at least $n^{c}$.
That is, wide-$\mathsf{TC}^{d}$ enforces that the size (and hence the width) of the circuit is large relative to its depth. In particular, for every wide-$\mathsf{TC}^{d}$ circuit family $\mathcal{C}$, there is a $c>0$ such that the circuit $C_{n}$ has serialization of size $\Omega(n^{c})$, and depth at most $c\log^{d}n$. Thus the depth of $C_{n}$ is at most $\log^{d}N$ where $N=n+\Omega(n^{c})$ is the overall size of the input $(x,\langle C_{n}\rangle)$ to the corresponding circuit evaluation problem. As we will show later, this allows the wide-$\mathsf{TC}^{d}$ circuit evaluation problem to be solved by a transformer (˜13.1) as well as by an $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ circuit family (Lemma˜14) using precisely $\log^{d}N$ iterations (manifested as loops of a transformer or $\mathsf{FO}$-uniform $\mathsf{TC}^{0}$ block, respectively), irrespective of the $\mathcal{C}$-dependent value of $c$.
Since it is a class of circuit families, wide-$\mathsf{TC}^{d}$ can be constrained by uniformity conditions in the natural way. With some abuse of notation, we will use wide-$\mathsf{TC}^{d}$ to refer to both the class of circuit families and the complexity class of language classes the circuit families recognize. Interestingly, this minimum size constraint imposed by wide-$\mathsf{TC}^{d}$ does not weaken it as a language class compared to $\mathsf{TC}^{d}$:
Proposition 4.
[…]
Proof.
Every wide-$\mathsf{TC}^{d}$ circuit family is, by definition, also a $\mathsf{TC}^{d}$ circuit family.
Conversely, suppose $L\in\mathsf{TC}^{d}$. Then there is a circuit family $\{C_{n}\}_{n=0}^{\infty}$ recognizing $L$, where the depth of $C_{n}$ is at most $c\log^{d}n$ for some $c$ and large enough $n$. Consider a modified circuit family $\{C^{\prime}_{n}\}_{n=0}^{\infty}$ where, for each $n$, $C^{\prime}_{n}$ is a copy of $C_{n}$ that, if needed, is padded with dummy gates so that it has size at least $n^{c}$. This modified circuit family also recognizes $L$ but belongs to wide-$\mathsf{TC}^{d}$, completing the proof. ∎
In fact, the above equality holds even for uniform variants of these classes:
Proposition 5.
[…]
Proof.
The proof follows that of Proposition˜4. Every $\mathsf{A}\textrm{-uniform}$ wide-$\mathsf{TC}^{d}$ circuit family is, by definition, also a $\mathsf{A}\textrm{-uniform}$ $\mathsf{TC}^{d}$ circuit family. Conversely, suppose $L\in\mathsf{A}$-uniform wide-$\mathsf{TC}^{d}$. Then, we have a a circuit family $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ recognizing $L$, where the depth of each $C_{n}$ is at most $c\log^{d}n$ for some $c$ and large enough $n$. By $\mathsf{A}$ uniformity, we have the connection language $L_{\mathcal{C}}\in\mathsf{A}$. We define a circuit family $\mathcal{C}^{\prime}$ that is $\mathcal{C}$ padded with dummy gates to make the number of gates at least $n^{c}$. The connection language $L_{\mathcal{C}^{\prime}}$ defaults to $L_{\mathcal{C}}$ for gates that exist in $\mathcal{C}$ and simply returns a dummy gate for other gate indices up to $n^{c}$. This additional logic can be implemented in $\mathsf{FO}$, so $L_{\mathcal{C}^{\prime}}\in\mathsf{A}\supseteq\mathsf{FO}$. Thus, the circuit family $\mathcal{C}^{\prime}$ recognizes $L$ and is $\mathsf{A}$-uniform. ∎
D.1 Hardness of Wide-$\mathsf{TC}^{d}$ Circuit Evaluation
@@
For $d\geq 0$, $\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}\textrm{-uniform}$ $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions.
@@
[…]
@@
For $d\geq 0$, wide-$\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}\textrm{-uniform}$ $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions.
@@
As in the proof of Lemma˜11, we are given $L$ such that there exists a log-space Turing machine $T_{L}$ that constructs $\{C_{n}\}_{n=0}^{\infty}$ recognizing $L$, where the depth of each $C_{n}$ is at most $c\log^{d}n$ for some $c$ and large enough $n$. We can create a modified log-space Turing machine $T^{\prime}_{L}$ that builds $\{C^{\prime}_{n}\}_{n=0}^{\infty}$ that still recognizes $L$ in depth $c\log^{d}n$ but is padded, if needed, with dummy gates so that it has size at least $n^{c}$: we do this by keeping a counter for the number of gates and wires produced and outputting dummy ones until $n^{c}$ is exceeded. We then follow the rest of the proof of Lemma˜11 with $T^{\prime}_{L}$ instead of $T_{L}$. ∎
D.2 Solving Wide-$\mathsf{TC}^{d}$ Circuit Evaluation with Transformers
@@
We adapt the proof of merrill-sabharwal-2023-parallelism, which sketches how log-depth (unlooped) transformers can implement the $\mathsf{TC}^{0}$ circuit evaluation problem. We construct a looped transformer that will “attempt” to evaluate every gate: if its arguments have already been computed, the gate will return $\{0,1\}$, and, if not, it will return undefined ($\bot$).
Let $i$ be a token index. We say the token $w_{i}$ is a gate token if it is $\texttt{X},\texttt{AND},\texttt{OR},\texttt{NOT},\texttt{MAJ}$. We will use gate token $i$ to store the value $v_{i}\in\{0,1,\bot\}$ for the gate it represents as a one-hot encoded vector. In the base case (embedding layer), we initialize $v_{i}=\bot$ for every gate token. Using a looped block of 2 layers, we will proceed in a way that propagates the computation of $v_{i}$ at later layers in terms of previously computed values.
@@
For $d\geq 0$, wide-$\mathsf{TC}^{d}$ circuit evaluation is in $\mathsf{mAHAT}^{d}_{0}$, and hence in $\mathsf{AHAT}^{d}_{1}$.
@@
We are given input $(x,\langle C_{n}\rangle)$, where $C_{n}$ comes from some wide-$\mathsf{TC}^{d}$ circuit family $\{C_{n}\}_{n=0}^{\infty}$ with depth at most $c\log^{d}n$ and size at least $n^{c}$, where $c$ is a constant specific to the circuit family. If we unroll the transformer from Lemma˜13 to depth $c\log^{d}n$, we can solve wide-$\mathsf{TC}^{d}$ circuit evaluation problem for large enough $n$ (w.l.o.g. we can solve small-$n$ examples via table lookup).
@@
Finally, we conclude via Proposition˜1 that this mixed-masked transformer can be converted to a corresponding causally masked padded transformer, placing the problem also in $\mathsf{AHAT}^{d}_{1}$. ∎
D.3 Solving Wide-$\mathsf{TC}^{d}$ Circuit Evaluation with $\mathsf{FO}$-Uniform Circuits
It is immediate that the looped transformer in ˜13.1 can be simulated in $\mathsf{L}$-uniform $\mathsf{TC}^{d}$. Moreover, since Lemma˜6 shows looped padded transformers can be simulated by $\mathsf{FO}$-uniform polylog-depth threshold circuit families, we obtain:
Lemma 14.
For $d\geq 0$, wide-$\mathsf{TC}^{d}$ circuit evaluation is in $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$.
Combined with Lemma˜12, this implies the following completeness results for wide-$\mathsf{TC}^{d}$ circuit evaluation:
Corollary 14.1.
For $d\geq 0$, wide-$\mathsf{TC}^{d}$ circuit evaluation is complete for $\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$ under $\mathsf{L}$ reductions.
Corollary 14.2.
For $d\geq 0$, wide-$\mathsf{TC}^{d}$ circuit evaluation is complete for $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ under $\mathsf{L}$ reductions.
Appendix E Uniformity Collapse for Circuit Classes
In general, for classes $\mathsf{A},\mathsf{B}$ such that $\mathsf{A}\subseteq\mathsf{B}$, $\mathsf{B}\textrm{-uniformity}$ often leads to larger classes of languages than $\mathsf{A}\textrm{-uniformity}$, as we have more resources to construct a circuit. However, as the following lemma shows, this is not always the case:
Proposition 6 (Uniformity Collapse).
Consider classes $\mathsf{A},\mathsf{B}$ of functions and a class $\mathsf{XC}$ of polynomial size circuit families such that:
@@ Appendix E Uniformity Collapse for Circuit Classes
    $\mathsf{XC}$ circuit evaluation is in $\mathsf{A}\textrm{-uniform}$ $\mathsf{XC}$;
  3. 3.
    $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$ circuit families are closed under fixed compositions (cf. Section˜B.1).
Then $\mathsf{B}\textrm{-uniformity}$ does not strengthen $\mathsf{A}\textrm{-uniformity}$, i.e., $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}=\mathsf{B}\textrm{-uniform}\ \mathsf{XC}$.
@@
Since $\mathsf{A}\subseteq\mathsf{B}$, we trivially have $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}\subseteq\mathsf{B}\textrm{-uniform}\ \mathsf{XC}$. The rest of the proof will focus on the other direction, showing that any language $L\in\mathsf{B}\textrm{-uniform}\ \mathsf{XC}$ is also in $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$.
By the definition of $\mathsf{B}\textrm{-uniform}\ \mathsf{XC}$, there exists a function $f\in\mathsf{B}$ that constructs an $\mathsf{XC}$ circuit family $\{C^{f}_{n}\}_{n\geq 0}$ that recognizes $L$. Specifically, $f(\texttt{1}^{n})=C^{f}_{n}$ where $C^{f}_{n}$ is an $\mathsf{XC}$ circuit that checks membership in $L$ over all strings $w$ of size $n$: for $w\in\{0,1\}^{n}$, $C^{f}_{n}(w)=1$ iff $w\in L$. The key insight is the following: while $\{C^{f}_{n}\}_{n\geq 0}$ is not necessarily an $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$ circuit family, we can build $C^{f}_{n}$ on demand using a different, $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$ circuit family $\{C^{g}_{n}\}_{n\geq 0}$, by leveraging the first precondition of the lemma, specifically that $B\subseteq\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$. We will then leverage the second precondition to construct another $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$ circuit family $\{C^{h}_{n,m}\}_{n,m\geq 0}$ that evaluates the built circuit $C^{f}_{n}$ on the input string $w$. We will then use the third precondition to compose these two circuit families in order to obtain the final $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$ circuit family $\{C^{L}_{n}\}_{n\geq 0}$ that recognizes $L$.
@@
Building $C^{f}_{n}$: Since $f\in\mathsf{B}$, by the first precondition of the lemma, $f$ is also in $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$. We can thus simulate $f$ using an $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$ circuit family. More concretely, there exists a function $g\in\mathsf{A}$ that, for any $n\geq 0$, maps input $\texttt{1}^{n}$ to an $\mathsf{XC}$ circuit $C^{g}_{n}$ that does what $f$ does, i.e., $C^{g}_{n}$ on input $\texttt{1}^{n}$ constructs the circuit $C^{f}_{n}$. Thus, we have $g(\texttt{1}^{n})=C^{g}_{n}$ and $C^{g}_{n}(\texttt{1}^{n})=C^{f}_{n}$.
Evaluating $C^{f}_{n}$ on $w$: By the second precondition of the proposition, there exists a function $h\in\mathsf{A}$ that, for any $n,m\geq 0$, maps input $(\texttt{1}^{n},\texttt{1}^{m})$ to an $\mathsf{XC}$ circuit $C^{h}_{n,m}$ that evaluates any $\mathsf{XC}$ circuit $C_{n}$ of size $m$ over any input string $w$ of size $n$, i.e., $h(\texttt{1}^{n},\texttt{1}^{m})=C^{h}_{n,m}$ and $C^{h}_{n,m}(C_{n},w)=C_{n}(w)$ for $\lvert w\rvert=n$.
Composing builder and evaluator circuits: Now we are ready to construct an $\mathsf{A}\textrm{-uniform}\ \mathsf{XC}$ circuit family $\{C^{L}_{n}\}_{n\geq 0}$ that recognizes $L$. This circuit family, on input $w$ of length $n$, computes the following composition of the circuit builder family $\{C^{g}_{n}\}_{n\geq 0}$ and the circuit evaluator family $\{C^{h}_{n,m}\}_{n,m\geq 0}$:
$C^{L}_{n}(w)=C^{h}_{n,m}(C^{g}_{n}(\texttt{1}^{n}),w)$ (1)
where $n=\lvert w\rvert$ and $m=\lvert C^{f}_{n}\rvert$. We first observe that this composite circuit does, in fact, recognize $L$. That’s because for any input string $w$ of length $n$:
$C^{L}_{n}(w)$ $=C^{h}_{n,m}(C^{g}_{n}(\texttt{1}^{n}),w)$
@@
$=C^{f}_{n}(w)$
which, by definition, is $1$ if and only if $w\in L$. Thus, $C_{L}$ recognizes $L$.
Lastly, by the third assumption of this proposition, the circuit family $\{C^{L}_{n}\}_{n\geq 0}$, being a fixed composition of $\mathsf{A}\textrm{-uniform}$ circuit families $\{C^{h}_{n,m}\}_{n,m\geq 0}$, $\{C^{g}_{n}\}_{n\geq 0}$, and the identify function (for constructing the second argument $w$ of $C^{h}_{n,m}$), is also $\mathsf{A}\textrm{-uniform}$. ∎